# Finte-size induced random switching of chimeras in a deterministic two-population Kuramoto-Sakaguchi model

### v1

@@ III Numerical results for the two-population Kuramoto-Sakaguchi model
To numerically integrate the two-population KS model (1)–(2) we prepare the initial conditions such that population $P_{1}$ is synchronised with initial phase $\theta_{1}^{(1)}=\theta_{2}^{(1)}=\dots=\theta_{N}^{(1)}=\psi_{1}=0$ and population $P_{2}$ is desynchronised with initial phases uniformly sampled from $\pi,\pi]$. We employ the MATLAB ode45 Runge-Kutta method to integrate (1)–(2) with relative and absolute tolerances set to 2.2 $2\,10^{-8}.$ To prevent an eventual formation of small clusters due to finite-size sample errors in the native frequencies Peter and Pikovsky (2018); Fialkowski et al. (2023), we draw the two sets of $N$ native frequencies $\omega^{(\alpha)}$ equiprobably. We show results for chimera states with $K=100$, $\kappa=60$ and $\lambda=\pi/2-0.075$. For these parameters the evolution equations (7)–(8) support a stable synchronised solution with constant mean-field variables.
To illustrate switching chimeras and how their existence depends on the number of oscillators $N$, we show in Figure 1 the temporal evolution of the order parameters $r_{1}$ and $r_{2}$ for $N$ ranging from $N=6$ to $N=128$. Switching between synchronised and desynchronised states is clearly seen. Switching events are not cyclic but appear randomly and the time between switches increases with $N$. We checked that for $N=21$ no chimera switching occurred for our initial condition when integrating for $10^{6}$ time units. This suggests that chimera switching is induced by finite-size effects. It is further seen that the synchronised population exhibits less variation in the order parameter for larger numbers of oscillators. Figure 2 shows the mean switching time $\bar{\tau}$, estimated from $1,000$ switching events, as a function of the system size $N$. The mean switching time $\bar{\tau}$ increases exponentially with the network size $N$. A best-fit line suggests $\bar{\tau}\sim\exp(0.889N)$. This implies that for all practical purposes chimera switching is a finite-size effect. In Figure 3 we show the empirical probability $P(\tau\geq\tau^{\star})$ of switching times $\tau$ which are greater than or equal to some $\tau^{\star}$ for fixed $N=12$, revealing an exponential distribution $P(\tau\geq\tau^{\star})\approx\exp(-\frac{1}{\bar{\tau}}\tau^{*})$ with $\bar{\tau}\approx 486.8$. This suggests that switching events occur independently and follow a Poisson process. We shall see that switching as well as the variance of the order parameter of the synchronised population is caused by an effective stochastic driving of the synchronised population by the desynchronised population.
@@
Figure 1: Order parameters $r_{1}$ (blue) and $r_{2}$ (orange) for the two-population KS model (1)–(2) exhibiting switching chimeras as a function of time for various numbers of oscillators $N$. Equation parameters are $K=100$, $\kappa=60$ and $\lambda=\frac{\pi}{2}-0.075$ and native frequencies $\omega$ are drawn equiprobably from a standard normal distribution $\mathcal{N}(0,1)$.
Figure 2: Mean switching time $\bar{\tau}$ as a function of the number of oscillators NN (blue circles). The orange line shows a best-fit exponential regression model with $\bar{\tau}=\exp(0.889N-4.15)$. Equation parameters are as in Figure 1.
Figure 3: Probability of switching times $P(\tau\geq\tau^{\star})$ as a function of $\tau^{\star}$ for a network of size $N=12$, estimated from $968$ switching events. Blue circles denote samples of the switching times. The orange line shows a best-fit regression model corresponding to $P(\tau\geq\tau^{\star})=\exp(-\tau^{\star}/\bar{\tau})$ with $\bar{\tau}\approx 486.8$. Equation parameters are as in Figure 1.
To investigate how such an effective stochastic dynamics is generated in the deterministic two-population KS model, we examine the statistical behaviour of $Z$, which captures the driving interaction term of the synchronised population $P_{1}$ with the desynchronised population $P_{2}$ (cf. (7)). Figure 4 shows the empirical distribution of the driving term $Z$ for a fixed number of oscillators $N=16$. To generate the empirical distribution we simulated (1)–(2) for $10,000$ time units, and determined $Z$ from the desynchronised population, employing the symmetry of the model, irrespective if that is population $P_{1}$ or $P_{2}$. The real and imaginary parts of $Z$ are well approximated by a Gaussian with an empirical mean of $\langle Z\rangle=0.6243+0.1007\,i$. The angular brackets $\langle\cdot\rangle$ denote a temporal average.
@@
is effectively a stochastic complex-valued Gaussian process. Hence, to obtain an effective stochastic dynamics capable of inducing the observed chimera switching and the approximation of $Z(t)$ by a Gaussian process, we require the number of oscillators to be sufficiently large for the central limit theorem to hold and sufficiently small so that the variance of the effective noise is nonnegligible. In Figure 6, we show how the empirical histograms of $\zeta$, obtained by simulations of (1)–(2) over $4,000$ time units, converge for increasing system size $N$. It is clearly seen that the fluctuations are not Gaussian for $N=8$, which is too small for the central limit theorem to work. For system sizes larger than $N=12$, the distribution of the components of $\zeta$ becomes approximately Gaussian, and indeed becomes more Gaussian as $N$ increases. It is seen that increasing the system size leads to a decrease in the variance of the Gaussian, with saturation occurring around $N=48$. For such systems with $N=48$, however, the phenomenon of switching is non-existent for all practical purposes. We hence limit the numerical study of switching to systems of $N\leq 16$.
@@
Recall that $Z$ encodes the effect of the desynchronised population onto the phases of the synchronised population, i.e. the second sum in the right-hand side of (1) using our convention that population $P_{1}$ is the synchronised one. This opens up the way to employ the stochastic model reduction framework developed in Yue and Gottwald (2024) to model the dynamics of the phases of the synchronised population. This will be achieved in Section V. To allow for an analytical approximation of the mean switching time we will need to determine the critical size of fluctuations that are sufficient to induce the synchronised population to switch. The asymptotic state of population $P_{1}$ sensitively depends on the initial condition of the desynchronised population $P_{2}$. Figure 7 shows the dependency of the average value of $\Delta r:=r_{2}-r_{1}=|Z|-r_{1}$ on the initial value $Z_{0}=Z(0)$ for a network with $N=512$ oscillators. Results were obtained by simulating the two-population KS model (1)–(2) for $10$ time units, for which the order parameters have equilibrated. We chose non-equilibrium initial conditions for population $P_{1}$ with $\theta_{1}^{(1)}=\theta_{2}^{(1)}=\dots=\theta_{N}^{(1)}$, and random initial conditions for population $P_{2}$ with $\theta^{(2)}_{j}$, $j=1,\dots,N$, drawn from the stationary density (14) consistent with a specified value of $Z_{0}$. We observe two scenarios: either $P_{1}$ is synchronised and $P_{2}$ is desynchronised (blue region), or $P_{1}$ is desynchronised and $P_{2}$ is synchronised (yellow region). The boundary separating these two states and its distance to the fixed point $\langle Z\rangle$ (red cross in Figure 7) determines the size of critical fluctuations of $Z$. The near circular boundary suggests that the critical fluctuation size is approximated by the average radius of the basin boundary which is estimated as $q^{\star}=0.386$. In the following Section we employ mean-field theory to determine an analytical expression for the critical size $q^{\star}$.
@@
Figure 5: Variance of the real (left) and imaginary (right) component of $Z$ as a function of the number of oscillators $N$. The orange lines show a power law best fit indicating a scaling with $N^{-1.13}$ and $N^{-1.10}$ for the real and imaginary parts, respectively. Equation parameters are as in Figure 1.
Figure 6: Empirical distributions of the real (a) and imaginary (b) part of $\zeta$ obtained from simulating (1)–(2) for $N=8$, $N=12$, $N=16$, $N=48$ and $N=96$, showing convergence with increasing system size $N$. The corresponding Kolmogorov–Smirnov statistics for the distributions of the real (blue) and imaginary (orange) components of $\zeta$ quantifies the convergence to a Gaussian with increasing NN (c). Equation parameters are as in Figure 1.
Figure 7: Time-averaged difference of the order parameters, $\langle\Delta r\rangle:=\langle r_{2}-r_{1}\rangle=\langle|Z|-r_{1}\rangle$, as a function of the initial condition $Z_{0}=Z(0)$ of the two-population KS model (1)–(2). The $N=512$ oscillators were initialised with $\theta_{j}^{(1)}(0)=0$ for all $j\in P_{1}$ and $\theta_{j}^{(2)}(0)$ were are drawn from the stationary distribution for non-entrained oscillators (14). The temporal mean is estimated during the time window $t\in[2.5,10]$. The thermodynamic equilibrium with $\langle Z\rangle=0.6230+0.0972i$ is shown as a red dot. The blue region denotes initial conditions which evolve into a state where $P_{1}$ is synchronised and $P_{2}$ is desynchronised, and the yellow region denotes initial conditions which evolve into states where $P_{1}$ is desynchronised and $P_{2}$ is synchronised. The black curve approximates the boundary separating these two states where we employed a Gaussian kernel to smooth the curve. The boundary has an average radius of $q^{\star}=0.386$. Equation parameters are as in Figure 1.
@@ IV Deterministic model reduction
Using the celebrated Ott-Antonsen ansatz valid in the thermodynamic limit, the two-population KS model (2)–(2) was shown to support stable chimeras, breathing chimeras, and saddle-node and Hopf bifurcations of chimera states Abrams et al. (2008); Laing (2009, 2012); Martens, Bick, and Panaggio (2016); Bick et al. (2020). We will follow Laing (2009) and derive here a low-dimensional dynamical model for the complex order parameters (3) in a perturbative approach. Using (7)–(8) we obtain the evolution equation for the order parameters
$\frac{d}{dt}z_{\alpha}=iz_{\alpha}^{(1,1)}+\frac{1}{2}(v_{\alpha}-v^{*}_{\alpha}z_{\alpha}^{(2,0)}),$ (20)
for the two populations with $\alpha=1,2$, where we write (8) as
$v_{1}$ $=(Kz_{1}+\kappa z_{2})e^{-i\lambda},$
$v_{2}$ $=(Kz_{2}+\kappa z_{1})e^{-i\lambda},$ (21)
and $z_{\alpha}^{(m,n)}$ is defined as
$z_{\alpha}^{(m,n)}=\frac{1}{N}\sum_{j=1}^{N}(\omega^{(\alpha)}_{j})^{n}\exp(im\theta^{(\alpha)}_{j})$, (22)
for $m,n\in\mathbb{N}$. Note that $z_{\alpha}^{(1,0)}=z_{\alpha}$.
Equation (20) is not a closed equation for the complex order parameters $z_{\alpha}$ but instead presents a closure problem whereby higher and higher orders of the order parameter are required to close the equation with
@@
We assume again without loss of generality that population $P_{1}$ is synchronised with $z_{1}=r_{1}e^{i\psi_{1}}$ and population $P_{2}$ is desynchronised with $z_{2}=r_{2}e^{i\psi_{2}}$. In order to close the equations for $z_{1,2}$ we truncate the system by employing the following two approximations
$z_{\alpha}^{(2,0)}=z_{\alpha}^{2},\qquad z_{\alpha}^{(1,1)}=0,$ (24)
for $\alpha=1,2$. We checked the validity of these approximations numerically (not shown), and provide a heuristic formal justification in Appendix A. Substituting the approximations (24) into (20) we obtain the closed Ott-Antonson equations Ott and Antonsen (2008); Laing (2012); Martens, Bick, and Panaggio (2016); Bick et al. (2020) for the dynamics of the complex order parameters with
@@
for $\alpha=1,2$. The closed equations (25) exhibit a Hopf bifurcation upon varying the intrapopulation coupling strength $K$ Abrams et al. (2008). Figure 8 shows the temporal evolution of the order parameters $r_{1,2}=|z_{1,2}|$ under the Ott-Antonson dynamics (25). The dynamics exhibits chimera states in the reduced model characterised by a synchronised population $P_{1}$ with $r_{1}=1$ and a desynchronised population $P_{2}$. The desynchronised population either has a stationary limiting state (with respect to the reference frame $\Omega$) with $r_{2}\approx 0.62$ or supports a stable limit cycle oscillating around a similar value. Such breathing chimera states are supported as well in the full two-population KS model (1)–(2). Our focus here is the case of a stationary fixed point. The stationary fixed point of the reduced Ott-Antonson equation (25) is Z=0.6096+0.0934iZ=0.6096+0.0934\,i, matching well the results of the full model (1)–(2) with $\langle Z\rangle=0.6243+0.1007\,i$ (cf. Figure 4).
Figure 8: Order parameters $r_{1}$ (blue) and $r_{2}$ (orange) as a function of time for the reduced deterministic mean-field dynamics (25) with $K=100$, $\kappa=60$ and $\lambda=\pi/2-0.075$ (left) and $K=90$, $\kappa=50$ and $\lambda=\pi/2-0.1$ (right). We used $z_{1}(0)=0.9$ and $z_{2}(0)=0.5$ to generate the chimera state in the reduced model.
The rapid relaxation of $r_{1}$ to its stable equilibrium at $r_{1}=1$ suggests to make the additional simplifying assumption of $r_{1}=1$, which allows us to express $Z$ as $Z=z_{2}/z_{1}=r_{2}e^{i\psi_{1}}$ (with $\Delta\psi=\psi_{1}-\psi_{2}$), which evolves according to
@@
The real scalar factor $\kappa\cos\lambda$ controls the time scale. Figure 9 shows the basin of attraction for the stationary desynchronised stationary fixed point and the desynchronised limit cycle from Figure 8 obtained from the reduced mean-field dynamics (IV). This compares well with the corresponding plots for the full two-population KS model (1)–(2) (cf. Figure 7).
Figure 9: Asymptotic value of the difference of the time-averaged order parameters, $\Delta r:=r_{2}-r_{1}=|Z|-r_{1}$, as a function of the initial condition $Z_{0}$ for the reduced deterministic mean-field model (IV). Left: Initial conditions in the blue region evolve to a fixed point with $\hat{Z}=0.6096+0.0934\,i$ (red dot) for $K=100$, $\kappa=60$ and $\lambda=\pi/2-0.075$. Initial conditions in the bluegreen region evolve into a fully synchronised state with $Z=1$. The boundary between the blue and green region is marked in light green. The red cross on the boundary marks a second unstable fixed point. We have overlaid the corresponding boundary curve (black) obtained from the full two-population KS model (1)–(2) for comparison (cf. Figure 7). Right: Initial conditions in the blue region evolve to a limit cycle (red loop) for $K=90$, $\kappa=50$ and $\lambda=\pi/2-0.1$. Initial conditions in the bluegreen region evolve into a fully synchronised state with $Z=1$. Equation parameters are as in Figure 8.
@@
$\frac{d}{dt}q$ $\approx Q(q)$
$:=\frac{\left(\int_{0}^{2\pi}\frac{\text{Re}(F(\hat{Z}+qe^{i\vartheta})e^{-i\vartheta})}{\text{Im}(F(\hat{Z}+qe^{i\vartheta})e^{-i\vartheta})}d\vartheta\right)}{\left(\int_{0}^{2\pi}\frac{1}{\text{Im}(F(\hat{Z}+qe^{i\vartheta})e^{-i\vartheta})}d\vartheta\right)},$ (27)
@@
Figure 10: Approximate temporal derivative of q=|ZZ^|q=|Z-$\hat${Z}|, $Q(q)$, as a function of $q$ obtained from numerically solving (IV) for $K=100$, $\kappa=60$, $\lambda=\frac{\pi}{2}-0.075$ (blue curve). A best-fit cubic approximation is shown as a red curve.

### v3

@@ III Numerical results for the two-population Kuramoto-Sakaguchi model
To numerically integrate the two-population KS model (1)–(2) we prepare the initial conditions such that population $P_{1}$ is synchronised with initial phase $\theta_{1}^{(1)}=\theta_{2}^{(1)}=\dots=\theta_{N}^{(1)}=\psi_{1}=0$ and population $P_{2}$ is desynchronised with initial phases uniformly sampled from $\pi,\pi]$. We employ MATLAB’s Runge-Kutta method ode45 to integrate (1)–(2) with relative and absolute tolerances of 2.2 $2\,10^{-8}.$ To prevent an eventual formation of small clusters due to finite-size sample errors in the native frequencies Peter and Pikovsky (2018); Fialkowski et al. (2023) and to ensure that both populations are exactly the same , we draw the two sets of $N$ native frequencies $\omega^{(\alpha)}$ equiprobably. We show results for chimera states with $K=100$, $\kappa=60$ and $\lambda=\pi/2-0.075$. For these parameters, the evolution equations (8)–(10) support a stable synchronised solution with constant mean-field variables.
To illustrate switching chimeras and how their existence depends on the number of oscillators $N$, we show in Figure 1 the temporal evolution of the order parameters $r_{1}$ and $r_{2}$ for $N$ ranging from $N=6$ to $N=128$. Switching between synchronised and desynchronised states is clearly seen. Switching events are not cyclic but appear randomly, and the time between switches increases with $N$. We checked that for $N=21$, no chimera switching occurred for our initial condition when integrating for $10^{6}$ time units. It is further seen that the synchronised population exhibits less variation in the order parameter for larger numbers of oscillators. Figure 2(a) shows the mean switching time $\bar{\tau}$, estimated from $1,000$ switching events, as a function of the system size $N$. The mean switching time $\bar{\tau}$ increases exponentially with the network size $N$. A best-fit line suggests $\bar{\tau}\sim\exp(0.889N)$. This suggests that chimera switching is a finite-size effect. In Figure 2(b) we show the empirical probability $P(\tau\geq t)$ of switching times $\tau$ which are greater than or equal to some $t$ for fixed $N=12$, revealing an exponential distribution $P(\tau\geq t)\approx\exp(-t/\bar{\tau})$ with $\bar{\tau}\approx 486.0$. This suggests that switching events occur independently and follow a Poisson process. We shall see that switching as well as the variance of the order parameter of the synchronised population is caused by an effective stochastic driving of the synchronised population by the desynchronised population.
@@
Figure 1: Order parameters $r_{1}$ (blue) and $r_{2}$ (red) for the two-population KS model (1)–(2) exhibiting switching chimeras as a function of time for various numbers of oscillators NN. Equation parameters are $K=100$, $\kappa=60$ and $\lambda=\frac{\pi}{2}-0.075$ and native frequencies $\omega$ are drawn equiprobably from a standard normal distribution $\mathcal{N}(0,1)$ such that both populations have the same set of native frequencies.
Figure 2: Statistics of the switching time. (a) Mean switching time $\bar{\tau}$ as a function of the number of oscillators NN (blue circles). The red line shows a best-fit exponential regression model with $\bar{\tau}=0.0158\times\exp(0.889N)$. (b): Empirical complementary cumulative distribution function $P(\tau\geq t)=\frac{1}{n}\sum_{i=1}^{n}\mathbf{1}_{\tau_{i}\geq t}$ for a network of size $N=12$, estimated from $n=1234$ switching events $\tau_{i}$ on a $\log$-scale (blue line). The red line shows a best-fit regression model corresponding to $P(\tau\geq t)=\exp(-t/\bar{\tau})$ with $\bar{\tau}\approx 486.0$. Equation parameters are as in Figure 1.
To investigate how such an effective stochastic dynamics is generated in the deterministic two-population KS model, we examine the statistical behaviour of $Z$, which captures the driving interaction term of the synchronised population $P_{1}$ with the desynchronised population $P_{2}$ (cf. (8)). Figure 3 shows the empirical distribution of the driving term $Z$ for $N=16$ oscillators. To generate the empirical distribution, we simulated (1)–(2) for $10,000$ time units, and determined $Z$ from the desynchronised population, employing the symmetry of the model, irrespective of whether that is population $P_{1}$ or $P_{2}$. The real and imaginary parts of $Z$ are well approximated by a Gaussian with an empirical mean of $\langle Z\rangle=0.6243+0.1007\,i$. The angular brackets $\langle\cdot\rangle$ denote a temporal average.
@@
is effectively a stochastic complex-valued Gaussian process. Hence, to obtain an effective stochastic dynamics capable of inducing the observed chimera switching and the approximation of $Z(t)$ by a Gaussian process, we require the number of oscillators to be sufficiently large for the central limit theorem to hold and sufficiently small so that the variance of the effective noise is non-negligible. In Figure 5 we show how the empirical histograms of the real and imaginary part of $\zeta$, obtained from simulations of (1)–(2) over $10,000$ time units, converge to a Gaussian for increasing system size $N$. We plot in Figure 5(a), (b) the empirical probability density function on a logarithmic scale after scaling the fluctuations by their empirical standard deviation and compare with a standard unit Gaussian distribution. It is clearly seen that the fluctuations are not Gaussian for $N=8$, which is too small for the central limit theorem to work. For larger system sizes, the distribution of the components of $\zeta$ becomes approximately Gaussian, and indeed becomes more Gaussian as $N$ increases. In particular, for $N=16$ the centre of the distribution of $\zeta$ is well approximated by a Gaussian but the tail for fluctuations deviating more than $2$ standard deviations from the mean departs from a Gaussian distribution. Note that this implies that roughly $94.5\%$ of the fluctuations are well described by a Gaussian distribution. We remark that increasing the system size leads to a decrease in the variance of the (scaled) fluctuations $\zeta$, with saturation occurring around $N=48$ (not shown). We further show in Figure 5(c) the Kolmogorov-Smirnov statistics $D^{*}$ quantifying the convergence to a Gaussian distribution with increasing system size MasseyFJ51 .
Note that the observed exponential increase of the switching time implies a mean switching time of approximately $5\times 10^{16}$ time units for $N=48$ (cf. Figure 2(a)). This implies that for all practical purposes, switching is not observable for $N=48$ in simulations, even for large integration times of $10^{6}$ time units. We hence limit the numerical study of switching to systems of $N\leq 16$.
@@
Recall that $Z$ encodes the effect of the desynchronised population onto the phases of the synchronised population, i.e. the second sum in the right-hand side of (1) using our convention that population $P_{1}$ is the synchronised one. This opens up the way to employ the stochastic model reduction framework developed in Yue and Gottwald (2024) to model the dynamics of the phases of the synchronised population. This will be achieved in Section V. To allow for an analytical approximation of the mean switching time, we will need to determine the critical size of fluctuations that are sufficient to induce the synchronised population to switch. The asymptotic state of population $P_{1}$ sensitively depends on the initial condition of the desynchronised population $P_{2}$. Figure 6 shows the dependency of the average value of $\Delta r:=r_{2}-r_{1}=|Z|-r_{1}$ on the initial value $Z_{0}=Z(0)$ for a network with $N=512$ oscillators. Results were obtained by simulating the two-population KS model (1)–(2) for $10$ time units, during which the order parameters have equilibrated. We chose non-equilibrium initial conditions for population $P_{1}$ with $\theta_{1}^{(1)}=\theta_{2}^{(1)}=\dots=\theta_{N}^{(1)}$, and random initial conditions for population $P_{2}$ with $\theta^{(2)}_{j}$, $j=1,\dots,N$, drawn from the stationary density (16) consistent with the initial value $Z_{0}$. We observe two scenarios: either $P_{1}$ is synchronised and $P_{2}$ is desynchronised (blue region), or $P_{1}$ is desynchronised and $P_{2}$ is synchronised (yellow region). The boundary separating these two states and its distance to the fixed point $\langle Z\rangle$ (red cross) determines the size of critical fluctuations of $Z$. The near circular boundary suggests that the critical fluctuation size is approximated by the average radius of the basin boundary, which is estimated as $q^{\star}=0.386$. In the following section, we employ mean-field theory to determine an analytical expression for the critical size $q^{\star}$.
@@
Figure 4: Variance of the real (left) and imaginary (right) component of $Z$ as a function of the number of oscillators NN. The red lines show a power law best fit indicating a scaling with $N^{-1.13}$ and $N^{-1.10}$ for the real and imaginary parts, respectively. Equation parameters are as in Figure 1.
Figure 5: Logarithm of the empirical distribution of the Z-score for the real (a) and imaginary (b) part of fluctuations $\zeta$ obtained from simulating (1)–(2) for $N=8$, $N=16$ and $N=96$ showing convergence with increasing system size $N$ to a standard normal distribution. The vertical lines delineate deviations with 22 standard deviations. (c): Kolmogorov–Smirnov statistics for the distributions of the real (blue) and imaginary (red) components of $\zeta$ quantifying the convergence to a Gaussian with increasing NN. Equation parameters are as in Figure 1.
Figure 6: Time-averaged difference of the order parameters, $\langle\Delta r\rangle:=\langle r_{2}-r_{1}\rangle=\langle|Z|-r_{1}\rangle$, as a function of the initial condition $Z_{0}=Z(0)$ of the two-population KS model (1)–(2). $N=512$ oscillators were initialised with $\theta_{j}^{(1)}(0)=0$ for all $j\in P_{1}$ and $\theta_{j}^{(2)}(0)$ were drawn from the stationary distribution for non-entrained oscillators (16). We used $30,000$ different initial conditions. The temporal mean was estimated during the time window $t\in[2.5,10]$. The thermodynamic equilibrium with $\langle Z\rangle=0.6230+0.0972i$ is shown as a red dot. The blue region denotes initial conditions that evolve into a state where $P_{1}$ is synchronised and $P_{2}$ is desynchronised, and the yellow region denotes initial conditions that evolve into states where $P_{1}$ is desynchronised and $P_{2}$ is synchronised. The black curve approximates the boundary separating these two states. The curve was obtained by first via ray casting from $\langle Z\rangle$ at $1,000$ different angles to locate points separating the two regions. The curve is then obtained by smoothing with a Gaussian kernel with a standard deviation of $0.162$ radians. The boundary has an average radius of $q^{\star}=0.386$. Equation parameters are as in Figure 1.
@@ IV Deterministic model reduction
Using the celebrated Ott-Antonsen ansatz valid in the thermodynamic limit, the two-population KS model (1)–(2) was shown to support stable chimeras, breathing chimeras, as well as saddle-node and Hopf bifurcations of chimera states Abrams et al. (2008); Laing (2009, 2012); Martens, Bick, and Panaggio (2016); Bick et al. (2020). We will follow Laing (2009) and derive a low-dimensional dynamical model for the complex order parameters (3) in a perturbative approach. Using (8)–(10) we obtain the evolution equation for the complex order parameters
$\frac{d}{dt}z_{\alpha}$ $=\frac{1}{N}\sum_{j\in P_{\alpha}}ie^{i\theta^{(\alpha)}_{j}}\left(\omega^{(\alpha)}_{j}+\text{Im}(v_{\alpha}e^{-i\theta^{(\alpha)}_{j}})\right)$
$=\frac{1}{N}\sum_{j\in P_{\alpha}}ie^{i\theta^{(\alpha)}_{j}}\left(\omega^{(\alpha)}_{j}+\frac{v_{\alpha}e^{-i\theta^{(\alpha)}_{j}}-v_{\alpha}^{*}e^{i\theta^{(\alpha)}_{j}}}{2i}\right)$
$=iz_{\alpha}^{(1,1)}+\frac{1}{2}(v_{\alpha}-v^{*}_{\alpha}z_{\alpha}^{(2,0)}),$ (23)
for the two populations with $\alpha=1,2$, $z_{\alpha}^{(m,n)}$ is defined as
$z_{\alpha}^{(m,n)}=\frac{1}{N}\sum_{j\in P_{\alpha}}(\omega^{(\alpha)}_{j})^{n}\exp(im\theta^{(\alpha)}_{j})$ (24)
for $m,n\in\mathbb{N}$, and where we again moved into a frame of reference of the synchronised cluster with mean phase $\psi_{1}$ and mean frequency $\Omega$, and where we have again assumed without loss of generality that population $P_{1}$ is synchronised with $z_{1}=r_{1}$ and $z_{2}=r_{2}e^{i(\psi_{2}-\psi_{1}})$. Note that $z_{\alpha}^{(1,0)}=z_{\alpha}$.
Equation (IV) is not a closed equation for the complex order parameters $z_{\alpha}$ but instead presents a closure problem whereby higher and higher orders of the order parameter are required to close the equation with
@@
In order to close the equations for $z_{1,2}$, we truncate the system by employing the following two approximations
$z_{\alpha}^{(2,0)}=z_{\alpha}^{2}\qquad{\rm{and}}\qquad z_{\alpha}^{(1,1)}=0$ (26)
for $\alpha=1,2$. We checked the validity of these approximations numerically (not shown) and provide a heuristic formal justification in Appendix A. Substituting the approximations (26) into (IV) we obtain the closed Ott-Antonsen equations Ott and Antonsen (2008); Laing (2012); Martens, Bick, and Panaggio (2016); Bick et al. (2020) for the dynamics of the complex order parameters with
@@
for $\alpha=1,2$. The closed equations (27) exhibit a Hopf bifurcation upon varying the intrapopulation coupling strength $K$ Abrams et al. (2008). Figure 7 shows the temporal evolution of the order parameters $r_{1,2}=|z_{1,2}|$ under the Ott-Antonsen dynamics (27). The dynamics exhibits chimera states in the reduced model characterised by a synchronised population $P_{1}$ with $r_{1}=1$ and a desynchronised population $P_{2}$. The desynchronised population either has a stationary limiting state (with respect to the reference frame $\Omega$) with $r_{2}\approx 0.62$ or supports a stable limit cycle oscillating around a similar value which are known as breathing chimera states in the full two-population KS model (1)–(2). Our focus here is on the case of a stationary fixed point. The stationary fixed point of the reduced Ott-Antonsen equation (27) is Z=0.6096+0.0934iZ=0.6096+0.0934\,i, matching well the results of the full model (1)–(2) with $\langle Z\rangle=0.6243+0.1007\,i$ (cf. Figure 3).
Figure 7: Order parameters $r_{1}$ (blue) and $r_{2}$ (red) as a function of time for the reduced deterministic mean-field dynamics (27) with $K=100$, $\kappa=60$ and $\lambda=\pi/2-0.075$ (left) and $K=90$, $\kappa=50$ and $\lambda=\pi/2-0.1$ (right). We used $z_{1}(0)=0.9$ and $z_{2}(0)=0.5$ to generate the chimera state in the reduced model.
The rapid relaxation of $r_{1}$ to its stable equilibrium at $r_{1}=1$ suggests to make the additional simplifying assumption of $r_{1}=1$, which allows us to express $Z$ as $Z=z_{2}/z_{1}=r_{2}e^{i\Delta\psi}$ (with $\Delta\psi=\psi_{1}-\psi_{2}$), which evolves according to
@@
The real scalar factor $\kappa\cos\lambda$ controls the time scale. Figure 8 shows the basin of attraction for the stationary desynchronised fixed point and the desynchronised limit cycle from Figure 7 obtained from the reduced mean-field dynamics (IV). This compares well with the corresponding plots for the full two-population KS model (1)–(2) (cf. Figure 6).
Figure 8: Asymptotic value of the difference of the time-averaged order parameters, $\Delta r:=r_{2}-r_{1}=|Z|-r_{1}$, as a function of the initial condition $Z_{0}$ for the reduced deterministic mean-field model (IV). Left: Initial conditions in the blue region evolve to a fixed point with $\hat{Z}=0.6096+0.0934\,i$ (red dot) for $K=100$, $\kappa=60$ and $\lambda=\pi/2-0.075$. Initial conditions in the teal region evolve into a fully synchronised state with $Z=1$. The boundary between the blue and teal regions is marked in green. The red cross on the boundary marks a second unstable fixed point. We have overlaid the corresponding boundary curve (black) obtained from the full two-population KS model (1)–(2) for comparison (cf. Figure 6). Right: Initial conditions in the blue region evolve to a limit cycle (red loop) for $K=90$, $\kappa=50$ and $\lambda=\pi/2-0.1$. Initial conditions in the teal region evolve into a fully synchronised state with $Z=1$. Equation parameters are as in Figure 7.
@@
$\frac{d}{dt}q$ $\approx Q(q):=\frac{\left(\int_{0}^{2\pi}\frac{\text{Re}(F(\hat{Z}+qe^{i\vartheta})e^{-i\vartheta})}{\text{Im}(F(\hat{Z}+qe^{i\vartheta})e^{-i\vartheta})}d\vartheta\right)}{\left(\int_{0}^{2\pi}\frac{1}{\text{Im}(F(\hat{Z}+qe^{i\vartheta})e^{-i\vartheta})}d\vartheta\right)},$ (29)
@@
Figure 9: Approximate temporal derivative of q=|ZZ^|q=|Z-$\hat${Z}|, $Q(q)$, as a function of $q$ obtained from numerically solving (29) for $K=100$, $\kappa=60$, $\lambda=\frac{\pi}{2}-0.075$ (blue curve). A best-fit cubic approximation is shown as a red curve.